A 50.0 mL solution of 0.150 M acetic acid (CH3
COOH) is titrated with 0.150 M NaOH.
What is the pH after 20.0 mL of base has been added? Ka
CH3
COOH = 1.8 x 10-5
1
Expert's answer
2020-11-18T13:43:21-0500

"\begin{aligned}
CH_3COOH + NaOH \to CH_3COONa + H_2O
\end{aligned}"

"\begin{aligned}
\textsf{No. of moles of }CH_3COOH &= Conc. × volume\\
&= 0.150M×\frac{50}{1000}\\
&= 0.0075 moles
\end{aligned}"

"\begin{aligned}
\textsf{No. of moles of }NaOH &= Conc. × volume\\
&= 0.150M×\frac{20}{1000}\\
&= 0.003moles
\end{aligned}"


After Neutralization, moles of excess acid = 0.0075 - 0.003 = 0.0045moles


"pH = pKa + log\dfrac{[A^-]}{[HA]}"


"\begin{aligned}
= -\log{1.8×10^{-5}} + \log\dfrac{0.003}{0.0045}
\end{aligned}"

"= 4.745 + 0.176"


"= 4.921"